OpenStudy (anonymous):

Can someone pleassssse help me with this problem? :) Solve for x. Enter only a number. 3 = log8 + 3logx HINT: If loga = 3, then a = 103.

OpenStudy (mrnood):

log8 is just a number so 3-log8 = 3logx or logx = (3-log8)/3 try it from there...

OpenStudy (akashdeepdeb):

a = 10^3 and not 103. 3 - log 8 = 3 * log x We know that 3 = log 10^3 So log 10^3 - log 8 = log 10^3 * log x Log rules tell us that log p - log q = log (p/q) Thus the resulting equation is log (10^3 / 8) = log 10^3 * log x Can you figure out the solution from here? :)

OpenStudy (anonymous):

I thinkkk so, thank you so much !

OpenStudy (anonymous):

\[3=\log 8++\log x^3=\log (8x^3),8x^3=10^3,(2x)^3=10^3\]

OpenStudy (anonymous):

2x=10 x=?